x - y = 0
2x + y = 12

Answer by richwmiller(17219)

solve by substitution or elimination
sub
x-y=0
x=y
2x+x=12
3x=12
x=4
y=4
eliminate
x-y=0
2x+y=12
add
3x=12
x=4
4-y=0
4=y